# Reduced dynamics for models of pattern formation

### v2

@@ Abstract.
Chemical and biochemical reactions can exhibit a wide range of complex behaviors, including multiple steady states, oscillatory patterns, and chaotic dynamics. These phenomena have captivated researchers for many decades. Notable examples of oscillating chemical systems include the Briggs–Rauscher, Belousov–Zhabotinskii, and Bray–Liebhafsky reactions, where periodic variations in concentration are often visualized through observable color changes. These systems are typically modeled by a set of partial differential equations coupled through nonlinear interactions.
Upon closer analysis, it appears that the dynamics of these chemical/biochemical reactions may be governed by only a finite number of spatial Fourier modes. We can also draw the same conclusion in fluid dynamics, where it has been shown that, over long periods, the fluid velocity is determined by a finite set of Fourier modes, referred to as determining modes. In this article, we introduce the concept of determining modes for a two-species chemical models, which covers models such as the Brusselator, the Gray-Scott model, and the Glycolysis model [1, 18]. We demonstrate that it is indeed sufficient to characterize the dynamic of the model using only a finite number of spatial Fourier modes.
@@
This work was supported by the Austrian Science Fund (FWF) 10.55776/P44681.
@@ 1. Introduction
The concept of determining modes becomes particularly relevant in this context. Originally introduced in the study of fluid dynamics and turbulence, the notion of determining modes is crucial for understanding complex phenomena like turbulence, which is characterized by chaotic and unpredictable fluid motion. Identifying and analyzing the underlying structures, or modes, within turbulent flows is essential for accurately predicting the behavior of fluid systems. Determining modes were introduced to identify the parameters that control turbulent flow; see, for example, the monograph by Foiaş [5, Chapter III] and the articles by Titi [10] and Foiaş and Temam [6]. This concept is practically significant. Modes are said “determining” if the high modes are asymptotically controlled by these lower modes. By identifying these determining modes, one can characterize the dominant patterns or structures within the flow. In the long run, a finite number of parameters — the coefficients of the determining modes — can effectively describe the entire system.
@@
(1.1) $\partial_{t}{u}(t)&=d_{1}\Delta u(t)+a_{1}u(t)+b_{1}v(t)-\gamma u(t)v^{2}(t)+g_{1}(t),&\mbox{ in }D,\\ \partial_{t}{v}(t)&=d_{2}\Delta v(t)+a_{2}u(t)+b_{2}v(t)+\gamma u(t)v^{2}(t)+g_{2}(t),&\mbox{ in }D,\\ u(0)&=u_{0}\geq 0,\;v(0)=v_{0}\geq 0,&\mbox{ in }D,\end{aligned}\right.$
@@
Concerning the solution theory, we can refer to [16], where the author discusses questions related to the global solution to a class of reaction-diffusion systems that satisfies two major properties: the quasi-positivity (P) and the mass-control (M). With these two properties valid, it is sometime possible to show the global existence of the classical solution. Moreover, the solution remains positive for all time. That applies, for instance, to the Brusselator model [7], which is a particular case of 1.1, where $a_{1}=a_{2}=0$; $b_{1}>0,b_{2}=b_{1}+1$; $\gamma=1$; $g_{1}=0,g_{2}>0$. The Brusselator satisfies the properties (P) + (M), and it was shown, see [8, Proposition 2], [17, Theorem 1, Page 140], that it has a unique global classical solution.
The system 1.1 fits also within the framework established by McGough and Riley in [14], where they develop a general class of reaction-diffusion problems characterized by telescoping nonlinearities, which arise in contexts such as chemical kinetics and population dynamics. These problems have special properties, which for the case of the two-species model 1.1 are given as follows:
@@
The goal of this paper is to demonstrate that the asymptotic behavior of the system of partial differential equations given by 1.1 can be effectively monitored by projecting the system onto a finite-dimensional space, known as determining modes. For a precise definition of the number of determining modes, we refer to Definition 2.3. In fluid dynamics, the number of determining modes is often expressed in terms of the Grashof number. This number quantifies the ratio of buoyancy forces to viscous forces in a fluid, making it an effective parameter for assessing the balance of forces within the system. For more details, see [5, Section II.13]. In our work, we will adopt the same approach, characterizing the number of determining modes using the Grashof number. To be more precise, we estimate the number of determining modes using a nondimensional constant, $\boldsymbol{\mathrm{Gr}}$, known as the Grashof number.
The rest of the paper is organized as follows. In Section 2, we collect background materials and necessary assumptions on the data and present the main result. In Section 3, we introduce some necessary Lemmata and propositions for the proof of the main statement. In Section 4, we prove that 1.1 admits a set of determining modes, see Definition 2.3.
@@ 2. Background materials and the main statement
In this section, we will introduce standard notations and recall results on the solution theory of systems like 1.1. Then, we present the definition of determining modes and the main result. The proof of the main result is postponed to Section 4.
[…]
@@
In the paper, $(u,v)$ is understood as the global smooth solution to 1.1, which existence is made precise in the following Lemmata, see also [14, Theorem 10].
@@
We assume that the properties ($P_{1})$-($P_{6})$ hold for the system 1.1. If $u_{0},v_{0}\in L^{2}(D)$ then for the solution to the system 1.1 it holds that $u(t),v(t)\geq 0$ and $u(t),v(t)\in C^{2}(D)$ for all $t\geq 0$.
More generally, the following estimates hold for $(u,v)$, see e.g. [14, Theorem 2].
@@ 2.1. Determining modes
To $(\tilde{g}_{1},\tilde{g}_{2})$ we associate the couple $(\tilde{u},\tilde{v})$ that satisfies for all $t\geq 0$ to the system
(2.2) $\partial_{t}{\tilde{u}}(t)&=d_{1}\Delta\tilde{u}(t)+a_{1}\tilde{u}(t)+b_{1}\tilde{v}(t)-\gamma\tilde{u}(t)\tilde{v}^{2}(t)+\tilde{g}_{1}(t),&\mbox{ in }D,\\ \partial_{t}{\tilde{v}}(t)&=d_{2}\Delta\tilde{v}(t)+a_{2}\tilde{u}(t)+b_{2}\tilde{v}(t)+\gamma\tilde{u}(t)\tilde{v}^{2}(t)+\tilde{g}_{2}(t),&\mbox{ in }D,\\ \tilde{u}(0)&=u_{0}\geq 0,\;\tilde{v}(0)=v_{0}\geq 0,&\mbox{ in }D,\end{aligned}\right.$
with zero Neumann boundary conditions on $\partial D$. We assume that the following properties holds for the system 2.2
@@
In what follows, $(\tilde{u},\tilde{v})$ is understood as the global smooth solution to 2.2. Lemma 2.1 and Lemma2.2 are also valid for the system 2.2 provided that the properties ($\tilde{P}_{1})$-($\tilde{P}_{6})$ hold.
Let $\{\psi_{m}:m\in\mathbb{N}\}$ be the complete set of eigenfunctions of the Laplacian and $\{\lambda_{m}:m\in\mathbb{N}\}$ be the associated eigenvalues such that $-\Delta\psi_{m}=\lambda_{m}\psi_{m},\;m\in\mathbb{N}$. Since $\Delta$ is selfadjoint, $\{\psi_{m}:m\in\mathbb{N}\}$ is an orthogonal basis of $L^{2}(D)$. Moreover, we suppose that
@@
Given some $M\in\mathbb{N}$, we define the finite subspace $\mathrm{H}_{M}\coloneqq\mbox{span}\big{\{}\psi_{m}:m\in\mathbb{N},\,m\leq M\big{\}}$ and denote the orthogonal projection of $L^{2}(D)$ onto $\mathrm{H}_{M}$ by $P_{M}$, the natural embedding by $I$, and define $Q_{M}$ by $Q_{M}\coloneqq I-P_{M}$. Given two indices $M$ and $N$ and the solution $(u,v)$ to 1.1, we define its Galerkin approximation by $(P_{M}u,P_{N}v)$.
@@
The first modes $H_{M}:=\big{\{}\psi_{m}:m\in\mathbb{N},\,m\leq M\big{\}}$ and $H_{N}:=\big{\{}\psi_{m}:m\in\mathbb{N},\,m\leq N\big{\}}$ are called determining modes for the solution $(u,v)$ to 1.1, if
@@ 2.2. The main statement
We set $\xi\coloneqq u-\tilde{u}$, $\eta\coloneqq v-\tilde{v}$, $h_{1}\coloneqq g_{1}-\tilde{g}_{1}$, and $h_{2}\coloneqq g_{2}-\tilde{g}_{2}$. We denote
$g_{1}^{*}\coloneqq\limsup_{t\to\infty}\|g_{1}(t)\|_{L^{2}}^{2},\quad g_{2}^{*}\coloneqq\limsup_{t\to\infty}\|g_{2}(t)\|_{L^{2}}^{2}.$
We set $F\coloneqq c_{1}+g_{1}^{*}+g_{2}^{*}+A_{2}^{2}+B_{2}$ and $d\coloneqq\min\{d_{1},d_{2}\}$, with $B_{2}\coloneqq c_{1}|D|^{1/2}$. We denote the generalized Grashof number that corresponds to the system 1.1 by
$\boldsymbol{\mathrm{Gr}}\coloneqq\frac{F}{d^{2}\lambda_{1}}.$
@@
We assume that the properties ($P_{1})$-($P_{6})$ hold for the system 1.1. If there exist $M,N\in\mathbb{N}$ such that
@@
then the first $(M,N)$ modes are determining for the solution $(u,v)$ to the system 1.1.
We postpone the proof of Theorem 2.4 to Section 4. In the first place, we show some preliminary estimates, notably Lemmas 3.1 , 3.2 and 3.3 that will be used as intermediate steps in proving Theorem 2.4.
@@ 3. Preliminaries
We will start with a Grownwall-type Lemma.
@@
[…]
@@
We deduce the results from the Gronwall lemma. See [4, Lemma 4.1] and [9, Lemma 4.1]. ∎
@@
We assume that the properties ($P_{1})$-($P_{6})$ (resp. ($\tilde{P}_{1})$-($\tilde{P}_{6})$) hold for the system 1.1 (resp. 2.2. We set
@@
We take the scalar product in $L^{2}(D)$ with $Q_{M}\xi$ and $Q_{N}\eta$ on both sides of 3.53.6, respectively. It follows that
@@
Next, we apply the Cauchy–Swcharz inequality, and for any $\varepsilon>0$, the Young inequality to get
@@
with $A_{4}\coloneqq|D|^{1/4}\sum_{j=1}^{2}(|a_{j}|+|b_{j}|)$. Next, we set
@@
We show that $\beta$ (resp. $\alpha$) satisfies to the hypothesis 3.2 (resp. 3.1) of Lemma 3.1 and conclude.
The hypothesis 3.2 of Lemma 3.1 on $\beta$ holds thanks to the asymptotic behavior 2.1, Lemma 2.2, and 3.4.
The hypothesis 3.1 of Lemma 3.1 on $\alpha$ holds as a consequence of 3.3. Indeed, observe that if
@@
Before we continue with 3.7, we show that
@@
with $A_{2}\coloneqq|D|^{1/2}\sum_{j=1}^{2}(|a_{j}|+|b_{j}|)$. By straightforward calculations, we get
@@
The proof of 3.8 is complete.
We can continue with 3.7, where by using the estimate 3.8, we get
@@
$\hskip 50.0pt-{4\Big{(}\gamma\|v(t)\|_{L^{4}}^{2}+\gamma\|\tilde{u}(t)\|^{2}_{L^{4}}+A_{2}\Big{)}\bigg{(}\frac{\|\nabla Q_{M}\xi(t)\|_{L^{2}}+\|\nabla Q_{N}\eta(t)\|_{L^{2}}}{\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}}\bigg{)}}-2\varepsilon$
@@
which by Lemma 2.2 provides 3.1. Now that every assumption of Lemma 3.1 holds, we can conclude that
@@
This completes the proof of Lemma 3.2. ∎
@@
We set $A_{2}\coloneqq|D|^{1/2}\sum_{j=1}^{2}(|a_{j}|+|b_{j}|)$. If there exist $T>0$ and $M,N\in\mathbb{N}$ such that
@@
then the first $(M,N)$ modes are determining for the solution $(u,v)$ to the system 1.1.
@@
×${}\times\bigg{(}\frac{1}{T}\int_{t}^{t+T}\big{(}H_{M,N}(\xi(s),\eta(s))\big{)}^{2}\,\mathrm{d}s\bigg{)}^{\frac{1}{2}}\geq\delta\times\bigg{(}\frac{1}{T}\int_{t}^{t+T}\big{(}H_{M,N}(\xi(s),\eta(s))\big{)}^{2}\,\mathrm{d}s\bigg{)}^{\frac{1}{2}}$
@@
If $T>0$ and $M,N\in\mathbb{N}$ are such that 3.9 holds, i.e. $\delta>0$, then it implies that
@@
in which case by Lemma 3.2, the first $(M,N)$ modes are determining for the system 1.1. ∎
4. Proof of Theorem 2.4
We split the proof of Theorem 2.4 into two Lemmata, then conclude. Let $(u,v)$ be any solution to the system 1.1.
@@
$\leq\int_{D}c_{1}(1-c_{2}u(t,x))u(t,x)\,\mathrm{d}x\leq c_{1}\big{(}\|u(t)\|_{L^{2}}-c_{2}\|u(t)\|_{L^{2}}^{2}\big{)}\leq B_{2}\|u(t)\|_{L^{2}},$
Where $B_{2}\coloneqq c_{1}|D|^{1/2}$.from which, on one hand, we obtain
@@
We recall, see [2, Proposition III.2.35], that
@@
$\frac{1}{2}\frac{\,\mathrm{d}\|w(t)\|_{L^{2}}^{2}}{\,\mathrm{d}t}+d_{1}\langle\nabla u(t),\nabla w(t)\rangle+d_{2}\langle\nabla v(t),\nabla w(t)\rangle=\langle(g_{1}+g_{2})(t),w(t)\rangle\leq\|(g_{1}+g_{2})(t)\|_{L^{2}}\|w(t)\|_{L^{2}}$
@@
which completes the proof of 4.2. ∎
Proof of Theorem 2.4
We fix $t\geq 0$ and $T>0$. To complete the proof of Theorem 2.4, we combine 4.14.2 to get
@@
Thus, by Lemma 3.3, the first $(M,N)$ modes are determining for the couple of solution $(u,v)$ to the system 1.1. This completes the proof of the main result Theorem 2.4. ∎
Acknowledgement
This work was supported by the Austrian Science Fund (FWF) 10.55776/P44681.

### v3

@@ Abstract.
The goal of this work is to analyze the long-term behavior of reaction-diffusion systems arising in two-species chemical models and to identify the minimal set of modes that determine their dynamics. The models considered include, as particular cases, the Brusselator, the Gray–Scott, and the Glycolysis models. These systems are described by coupled reaction-diffusion equations and admit a finite-dimensional representation based on a limited number of spatial Fourier modes that capture their essential reduced dynamics. The concept of determining modes, introduced in this context, is closely related to other approaches that seek finite-dimensional representations of infinite-dimensional dynamics, such as the Proper Orthogonal Decomposition and the construction of Approximate Inertial Manifolds. We prove that the dynamics of the system can be completely characterized by a finite number of low modes, since all higher modes are asymptotically determined by them, thus providing an analytical foundation for reduced dynamics in models of pattern formation.
@@
This work was supported by the Austrian Science Fund (FWF) 10.55776/P44681. The second author was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 317210226 - SFB 1283.
@@ 1. Introduction
The concept of determining modes becomes particularly relevant in this context. Originally introduced in the study of fluid dynamics and turbulence, the notion of determining modes is crucial for understanding complex phenomena like turbulence, which is characterized by chaotic and unpredictable fluid motion. Identifying and analyzing the underlying structures, or modes, within turbulent flows is essential for accurately predicting the behavior of fluid systems. Determining modes were introduced to identify the parameters that control turbulent flow; see, for example, the monograph by Foiaş [7, Chapter III] and the articles by Titi [13] and Foiaş and Temam [5]. Modes are said “determining” if the high modes are asymptotically controlled by these lower modes. By identifying these determining modes, one can characterize the dominant patterns or structures within the flow. In the long run, a finite number of parameters, the coefficients of the determining modes, can effectively describe the entire system.
@@
(1.1) $\left\{\begin{aligned} \partial_{t}{u}(t)&=d_{1}\Delta u(t)+a_{1}u(t)+b_{1}v(t)-\gamma u(t)v^{2}(t)+g_{1}(t),&\mbox{ in }D,\\ \partial_{t}{v}(t)&=d_{2}\Delta v(t)+a_{2}u(t)+b_{2}v(t)+\gamma u(t)v^{2}(t)+g_{2}(t),&\mbox{ in }D,\\ u(0)&=u_{0}\geq 0,\;v(0)=v_{0}\geq 0,&\mbox{ in }D,\end{aligned}\right$.
@@
Concerning the solution theory, we can refer to [20], where the author discusses questions related to the global solution to a class of reaction-diffusion systems that satisfies two major properties: the quasi-positivity (P) and the mass-control (M). With these two properties valid, it is sometime possible to show the global existence of the classical solution. Moreover, the solution remains positive for all time. That applies, for instance, to the Brusselator model [10], which is a particular case of (1.1), where $a_{1}=a_{2}=0$; $b_{1}>0,b_{2}=b_{1}+1$; $\gamma=1$; $g_{1}=0,g_{2}>0$. The Brusselator satisfies the properties (P) + (M), and it was shown, see [11, Proposition 2], [22, Theorem 1, Page 140], that it has a unique global classical solution. This result applies also to classical pattern-forming systems such as the Gray–Scott, and Glycolysis models [1, 23].
The system (1.1) fits also within the framework established by McGough and Riley in [18], where they develop a general class of reaction-diffusion problems characterized by telescoping nonlinearities, which arise in contexts such as chemical kinetics and population dynamics. These problems have special properties, which for the case of the two-species model (1.1) are given as follows:
@@
The goal of this paper is to demonstrate that the asymptotic behavior of the system of partial differential equations given by (1.1) can be effectively monitored by projecting the system onto a finite-dimensional space, known as determining modes. For a precise definition of the number of determining modes, we refer to Definition 2.1.
The analysis presented here is closely connected to a recent work on data assimilation for the Gray–Scott model [21], where synchronization of a nudged solution with the true dynamics was achieved from coarse cell-averaged measurements. In that framework, feedback is applied only to low finite-volume modes through an interpolation operator $\mathcal{I}_{H}$. The present determining modes result provides the analytical foundation for such synchronization: Theorem˜2.3 shows that, once the low modes, corresponding to the observed scales, are controlled, the remaining unobserved high modes decay asymptotically. In this sense, the determining subspace in the current work coincides with the set of spatial modes on which feedback or observation must act to ensure recovery of the full Gray–Scott dynamics. The two results, therefore, describe complementary aspects of the same phenomenon: one from the perspective of dynamical systems theory, the other from the perspective of control and data assimilation.
Beyond data assimilation, the present work is also connected to other approaches that seek finite-dimensional representations of infinite-dimensional dynamics. Among the most widely used are the Proper Orthogonal Decomposition (POD) and the construction of Approximate Inertial Manifolds (AIM). Both methods aim to identify low-dimensional structures that capture the long-term dynamics of dissipative partial differential equations.
In POD, one seeks a finite set of empirical orthogonal modes that optimally represent the energy content of the system; see, for instance, the monographs by Berkooz, Holmes, and Lumley [2], and the later developments summarized in [7]. The construction of AIMs, introduced by Foiaş, Sell, and Temam [9], provides a rigorous framework in which high modes are expressed as Lipschitz functions of the low modes, thereby yielding a finite-dimensional approximation of the global attractor. In contrast, the determining modes approach establishes explicit analytical conditions ensuring that a finite set of modes uniquely determines the full dynamic.
Related ideas have been further developed in the context of determining nodes, determining volume elements, and determining forms, see, e.g., Foiaş, Jolly, Kravchenko, and Titi [6]; Foiaş and Temam [5]; Jones and Titi [12, 13]; and the work of Kaper, Wang, and Wang on nonlinear Ginzburg–Landau equations [15]. These frameworks share the same objective: to quantify the finite number of degrees of freedom needed to describe the long-time dynamics of dissipative systems. Our present results extend these ideas to reaction-diffusion systems relevant to pattern formation, including the Gray–Scott, Brusselator, and Glycolysis models.
The rest of the paper is organized as follows. In Section˜2, we collect background materials and necessary assumptions on the system and present the main result. In Section˜3, we introduce some necessary lemmas and propositions that serve as intermediate steps in the proof of the main result. In Section˜4, we prove that the reaction-diffusion system (1.1) admits a set of determining modes, as defined in Definition˜2.1. Finally, in Section˜5, we discuss the implications of the main result for Turing-pattern models, including its connection to data assimilation for the Gray–Scott system.
@@ 2. Background materials and the main statement
In this section, we will introduce standard notations and recall results on the solution theory of systems like (1.1). Then, we present the definition of determining modes and the main result. The proof of the main result is postponed to Section˜4.
By $L^{2}\coloneqq L^{2}(D)$, we denote the usual Lebesgue space of squared integrable functions. It is endowed with the scalar product $\langle\,\cdot\,$$\cdot\,\rangle$. For $1\leq p\leq\infty$, we denote the usual Lebesgue spaces by $L^{p}$, which are endowed with the standard norms denoted by
@@
In the paper, $(u,v)$ is understood as the global smooth solution to (1.1), which existence is made precise in the following lemmas, see also [18, Theorem 10].
@@
We assume that the properties ($P_{1})$-($P_{6})$ hold for the system (1.1). If $u_{0},v_{0}\in L^{2}(D)$ then for the solution to the system (1.1) it holds that $u(t),v(t)\geq 0$ and $u(t),v(t)\in C^{2}(D)$ for all $t\geq 0$.
More generally, the following estimates hold for $(u,v)$, see e.g. [18, Theorem 2].
@@ 2.1. Determining modes
[…]
(2.2) $\left\{\begin{aligned} \partial_{t}{\tilde{u}}(t)&=d_{1}\Delta\tilde{u}(t)+a_{1}\tilde{u}(t)+b_{1}\tilde{v}(t)-\gamma\tilde{u}(t)\tilde{v}^{2}(t)+\tilde{g}_{1}(t),&\mbox{ in }D,\\ \partial_{t}{\tilde{v}}(t)&=d_{2}\Delta\tilde{v}(t)+a_{2}\tilde{u}(t)+b_{2}\tilde{v}(t)+\gamma\tilde{u}(t)\tilde{v}^{2}(t)+\tilde{g}_{2}(t),&\mbox{ in }D,\\ \tilde{u}(0)&=u_{0}\geq 0,\;\tilde{v}(0)=v_{0}\geq 0,&\mbox{ in }D,\end{aligned}\right$.
with zero Neumann boundary conditions on $\partial D$. We assume that the following properties holds for the system (2.2)
@@
[…]
Let $\{\psi_{m}:m\in\mathbb{N}\}$ be the complete set of eigenfunctions of the Laplacian and $\{\lambda_{m}:m\in\mathbb{N}\}$ be the associated eigenvalues such that $\Delta\psi_{m}=\lambda_{m}\psi_{m},\;m\in\mathbb{N}$. Since $\Delta$ is selfadjoint, $\{\psi_{m}:m\in\mathbb{N}\}$ is an orthogonal basis of $L^{2}(D)$. Moreover, we suppose that
@@
Given some $M\in\mathbb{N}$, we define the finite subspace $\mathrm{H}_{M}\coloneqq\mbox{span}\big\{\psi_{m}:m\in\mathbb{N},\,m\leq M\big\}$ and denote the orthogonal projection of $L^{2}(D)$ onto $\mathrm{H}_{M}$ by $P_{M}$, the natural embedding by $I$, and define $Q_{M}$ by $Q_{M}\coloneqq I-P_{M}$. Given two indices $M$ and $N$ and the solution $(u,v)$ to (1.1), we define its Galerkin approximation by $(P_{M}u,P_{N}v)$.
@@
The first modes $H_{M}:=\big\{\psi_{m}:m\in\mathbb{N},\,m\leq M\big\}$ and $H_{N}:=\big\{\psi_{m}:m\in\mathbb{N},\,m\leq N\big\}$ are called determining modes for the solution $(u,v)$ to (1.1), if
@@ 2.2. The main statement
We introduce the notation
$\xi\coloneqq u-\tilde{u},\qquad\eta\coloneqq v-\tilde{v},\qquad h_{1}\coloneqq g_{1}-\tilde{g}_{1},\qquad h_{2}\coloneqq g_{2}-\tilde{g}_{2}.$
We set
$g_{1}^{*}\coloneqq\limsup_{t\to\infty}\|g_{1}(t)\|_{L^{2}}^{2},\qquad g_{2}^{*}\coloneqq\limsup_{t\to\infty}\|g_{2}(t)\|_{L^{2}}^{2}.$
Let $c_{D}$ denote the Lebesgue measure of the domain $D$, and define
$A_{2}\coloneqq c_{D}^{1/2}\sum_{j=1}^{2}(|a_{j}|+|b_{j}|),\qquad B_{2}\coloneqq c_{1}c_{D}^{1/2}.$
We further set
$d\coloneqq\min\{d_{1},d_{2}\},\qquad F\coloneqq c_{1}+g_{1}^{*}+g_{2}^{*}+A_{2}^{2}+B_{2},\qquad\boldsymbol{\mathrm{Gr}}\coloneqq\frac{F}{d^{2}\lambda_{1}}.$
@@
We assume that the properties ($P_{1})$-($P_{6})$ hold for the system (1.1). If there exist $M,N\in\mathbb{N}$ such that
@@
then the first $(M,N)$ modes are determining for the solution $(u,v)$ to the system (1.1).
We postpone the proof of Theorem˜2.3 to Section˜4. In the first place, we show some preliminary estimates, notably Sections˜3 , 3 and 3 that will be used as intermediate steps in proving Theorem˜2.3.
@@ 3. Preliminaries
We will start with a Gronwall-type Lemma.
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}X(t)+\alpha(t)X(t)\leq\beta(t),\;\mbox{a.e.\ on }(0,\infty).$
@@
We deduce the results by the Gronwall lemma. See [8, Lemma 4.1] and [12, Lemma 4.1]. ∎
@@
We assume that the properties ($P_{1})$-($P_{6})$ (resp. ($\tilde{P}_{1})$-($\tilde{P}_{6})$) hold for the system (1.1) (resp. (2.2). We set
@@
We take the scalar product in $L^{2}(D)$ with $Q_{M}\xi$ and $Q_{N}\eta$ on both sides of (3.5)-(3.6), respectively. It follows that
@@
Next, we apply the Cauchy–Schwarz inequality, and for any $\varepsilon>0$, the Young inequality to get
@@
with $A_{4}\coloneqq c_{D}^{1/4}\sum_{j=1}^{2}(|a_{j}|+|b_{j}|)$. Next, we set
@@
We show that $\beta$ (resp. $\alpha$) satisfies to the hypothesis (3.2) (resp. (3.1)) of Section˜3 and conclude.
The hypothesis (3.2) of Section˜3 on $\beta$ holds by the asymptotic behavior (2.1), Section˜2, and (3.4).
The hypothesis (3.1) of Section˜3 on $\alpha$ holds as a consequence of (3.3). Indeed, observe that if
@@
Before we continue with the proof of (3.7), we show that
@@
with $A_{2}\coloneqq c_{D}^{1/2}\sum_{j=1}^{2}(|a_{j}|+|b_{j}|)$. By straightforward calculations, we get
@@
The proof of (3.8) is complete.
We can continue with the proof of (3.7), where by using the estimate (3.8), we get
@@
$\hskip 25.0pt-{4\Big(\gamma\|v(t)\|_{L^{4}}^{2}+\gamma\|\tilde{u}(t)\|^{2}_{L^{4}}+A_{2}\Big)\bigg(\frac{\|\nabla Q_{M}\xi(t)\|_{L^{2}}+\|\nabla Q_{N}\eta(t)\|_{L^{2}}}{\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}}\bigg)}-2\varepsilon$
@@
which by Section˜2 provides (3.1). Now that every assumption of Section˜3 holds, we can conclude that
@@
This completes the proof of Section˜3. ∎
@@
We recall that $A_{2}\coloneqq c_{D}^{1/2}\sum_{j=1}^{2}(|a_{j}|+|b_{j}|)$. If there exist $T>0$ and $M,N\in\mathbb{N}$ such that
@@
then the first $(M,N)$ modes are determining for the solution $(u,v)$ to the system (1.1).
@@
${}\times\bigg(\frac{1}{T}\int_{t}^{t+T}\big(H_{M,N}(\xi(s),\eta(s))\big)^{2}\,\mathrm{d}s\bigg)^{\frac{1}{2}}\geq\delta\times\bigg(\frac{1}{T}\int_{t}^{t+T}\big(H_{M,N}(\xi(s),\eta(s))\big)^{2}\,\mathrm{d}s\bigg)^{\frac{1}{2}}$
@@
If $T>0$ and $M,N\in\mathbb{N}$ are such that (3.9) holds, i.e. $\delta>0$, then it implies that
@@
in which case by Section˜3, the first $(M,N)$ modes are determining for the system (1.1). ∎
4. Proof of Theorem˜2.3
We split the proof of Theorem˜2.3 into two lemmas, then conclude. Let $(u,v)$ be any solution to the system (1.1).
@@
$\leq\int_{D}c_{1}(1-c_{2}u(t,x))u(t,x)\,\mathrm{d}x\leq c_{1}\big(\|u(t)\|_{L^{1}}-c_{2}\|u(t)\|_{L^{2}}^{2}\big)\leq B_{2}\|u(t)\|_{L^{2}},$
where $B_{2}\coloneqq c_{1}c_{D}^{1/2}$, from which, on one hand, we obtain
@@
We recall, see [3, Proposition III.2.35], that
@@
$\frac{1}{2}\frac{\,\mathrm{d}\|w(t)\|_{L^{2}}^{2}}{\,\mathrm{d}t}+d_{1}\langle\nabla u(t),\nabla w(t)\rangle+d_{2}\langle\nabla v(t),\nabla w(t)\rangle$
$=\langle(g_{1}+g_{2})(t),w(t)\rangle\leq\|(g_{1}+g_{2})(t)\|_{L^{2}}\|w(t)\|_{L^{2}}$
@@
which completes the proof of (4.2). ∎
Proof of Theorem˜2.3
We fix $t\geq 0$ and $T>0$. To complete the proof of Theorem˜2.3, we combine (4.1)-(4.2) to get
@@
Thus, by Section˜3, the first $(M,N)$ modes are determining for the couple of solution $(u,v)$ to the system (1.1). This completes the proof of the main result Theorem˜2.3. ∎
5. Discussion
The recent work “A discrete data assimilation algorithm for the reconstruction of Gray–Scott dynamics” [21] demonstrates that one can reconstruct the full Gray–Scott solution from coarse, cell-averaged data via the Azouani–Olson–Titi (AOT) data assimilation algorithm, as given by
[…]
How many modes of the Gray–Scott system or another model of pattern formation are sufficient to determine the full dynamics?
The determining modes framework answers this question analytically at the continuous PDE level, without referring to discrete data or to a specific data assimilation algorithm.
The structure of synchronization in the AOT system is, in fact, mathematically analogous to that of the determining modes property proved here in Theorem˜2.3. Indeed, the determining modes result can be expressed as
$\lim_{t\to\infty}\big(\|P_{M}(u-\tilde{u})\|_{L^{4}}+\|P_{N}(v-\tilde{v})\|_{L^{4}}\big)=0\quad\mbox{implies}\quad\lim_{t\to\infty}\big(\|Q_{M}(u-\tilde{u})\|_{L^{2}}+\|Q_{N}(v-\tilde{v})\|_{L^{2}}\big)=0.$
In both settings, the low modes (or coarse observations) act as determining observables. The high modes (fine details) are asymptotically determined to those determining observables. The constants and inequalities connecting the feedback parameters $(\mu_{u},\mu_{v},H)$ in the data assimilation framework correspond analytically to the determining mode inequality
${\lambda_{M+1}+\lambda_{N+1}}>\lambda_{1}(2\gamma^{2}d^{2}\boldsymbol{\mathrm{Gr}}^{4}+\boldsymbol{\mathrm{Gr}}).$
Hence, the main result Theorem˜2.3 of the present paper provides the analytical foundation for why data assimilation via coarse measurements can succeed: the feedback in the AOT algorithm acts precisely on the determining subspace that governs the full system dynamics. This correspondence highlights the central idea of the present work, notably that complex reaction-diffusion systems, such as the Gray–Scott or Brusselator models, admit a finite-dimensional description capturing their essential reduced dynamics.
Declarations
Conflicts of Interest
The authors have no competing interests to declare that are relevant to the content of this article.